# ThinkTank: A Benchmark for Cognitive Intelligence in Pre-trained Language Models

### v1

@@ 3.4 Fluid reasoning
Humans can logically parse information and detect patterns in novel situations without having to rely on prior experiences or learned information. This ability is called Fluid Reasoning cattell1963theory.
@@
Amber liu2023llm360
Falcon falcon40b
Starling starling2023
Llama touvron2023llama
Mistral jiang2023mistral
Qwen qwen
@@
biderman2023pythia
Gemini geminiteam2023gemini
GPT-3.5-Turbo openai
GPT 4 openai2023gpt4
@@
We focus on the original measure of fluid reasoning, the Ravens Progressive Matrices (RPM) test raven2003raven. An example Ravens-like problem is given in Figures 1D and 3. An RPM item consists of a 3x3 matrix of cells with one empty cell. Participants must induce the underlying, abstract patterns that hold across the rows and columns of the matrix, and apply these to infer the image in the empty cell from a given set of options. These images vary in visual attributes like shape and color, along with more abstract qualities. The RPM is the standard measure of fluid reasoning snow1984topography.
@@
Given the visual nature of the RPM, hu2021stratified; hu2023context mapped the Raven-10000 dataset to a textual format to facilitate the testing of PLMs. The mapping involves reformulating visual elements into text-based numerical tuples to form the I-Raven dataset, representing attributes like shape, size, and color textually, as illustrated in Figure 3. We use their approach with a focus on the “Center Single Alignment” sub-task, which features a single shape per matrix cell. We differ from their work by evaluating a broader set of models.
Table 3: Performance of Pre-trained Language Models on ThinkTank. Distance Effect: Averaged $R^{2}$ values of different LLMs when fitting a linear function on the cosine-similarity vs. distance plot. Ratio Effect: Averaged $R^{2}$ values of different LLMs when fitting a negative exponential function on the cosine-similarity vs. ratio plot. Note: Each value is averaged across all three input types and all model layers to produce one generalizable score. Latent Rep: Average Spearman’s Correlation when using the cosine similarity and latent representation-based approach (Note: * refers to the prompting approaches for select models which are gated by APIs and not the latent representation-based approach), Zero-Shot: Average Spearman’s Correlation when using the zero-shot surprisal values, BLiMP: The Benchmark of Linguistic Minimal Pairs for English, RPM: Raven’s Progressive Matrices
@@
Operationalization: We determine the model’s preferred answer for a problem by comparing the surprisal values of the whole sequence (instruction, question, candidate tuple) for each of the candidate options, i.e. the probability of each completed digit representation of a matrix. For the example given in Figure 3, this would be checking the probability of this sequence (summation of token probabilities) with the correct answer (3, 0.6, 0.8) to the other candidates. A comprehensive list of the prompts used in this paper is given in Appendix B.4.
@@ 4 Models under consideration
We evaluate a wide range of language model families, shown in Table 2, on ThinkTank. These models are selected based on the following criteria:
  • Public availability: Open-source models allow us to perform a thorough analysis by accessing the latent representation and the token probability during generation. Although most of the models in this study are publicly available and open-source, we use three state-of-art commercial PLMs that are gated behind API calls; GPT-3.5-Turbo (pointing to gpt-3.5-turbo-0613 on the OpenAI platform), GPT-4 (pointing to gpt-4-1106-preview on the OpenAI platform), and Gemini (also referred to as Gemini-1-Pro at the time of writing). The GPT-$x$ model APIs provide token probabilities of the response, allowing us to calculate surprisal, while Gemini does not.
@@
  • Availability of intermediate training checkpoints: This allows us to evaluate the effects of pre-training on the model outputs. Together, the availability of multiple model sizes and intermediate training checkpoints allow us to best evaluate the developmental alignment of PLMs. Amber and Pythia’s family of models have available intermediate training checkpoints. While Amber has 360 intermediate training checkpoints, the checkpoints are at 4 Billion tokens each and are not at the required granularity.
Pythia Family of models: Pythia biderman2023pythia is one of the first open-source projects with the goal of scientific and transparent model development. It has 8 model sizes with parameters ranging from 70 Million to 12 Billion, with each model trained on 286 Billion tokens. The models in the suite are equivalent (in size) to popular decoder architectures like GPT-Neo-(125M, 1.3B, 2.7B) and OPT-(125M, 350M, 1.3B, 2.7B, 6.7B), but with the added benefits of training on a known de-duplicated corpus gao2020pile, the same training order for each model size, and having 154 intermediate checkpoints to study the learning trajectories of PLMs. Thus, the Pythia suite of models is ideal for studying the psychometric and developmental alignment of PLMs to humans.
Note: All open-source models are obtained from Huggingface huggingface, while the gated models are obtained from their respective platforms through API calls. For each model in the Pythia suite, the following intermediate checkpoints are available: [1, 2, 4, 8, … 512; 1000, 2000, 3000 … 143000 (exponential increase in checkpoint number till 512th checkpoint and subsequent progression of 1000 steps till the last checkpoint)], with each checkpoint representing 2 Million tokens seen.
@@ 5 Cognitive and developmental alignment of PLMs
ThinkTank enables comprehensive evaluation of a variety of PLMs on their cognitive alignment to humans across four domains of psychometric intelligence: numeric abilities, linguistic abilities, concept understanding, and fluid reasoning. Table 3.4 highlights the key results of this evaluation. Note: for the evaluation of conceptual understanding in PLMs, we only report the results for the zero-shot surprisal values and latent representations. This is because we see similar results for zero-shot and few-shot surprisal value-based methods (see comprehensive results in Appendix B.3).
@@
  • Linguistic abilities: The accuracy of the PLMs on the BLiMP linguistic acceptability tasks improves upon increasing the number of parameters. Furthermore, we find that all PLMs are substantially more accurate on morphological tasks over syntactic and semantic tasks (Accuracy: semantic $<$ syntax <<much-less-than$<$ < morphology; see Appendix Table 5).
  • Concept understanding: Prompting methods in commercial models perform substantially better than both methods – closeness judgment and surprisal values – on all open-source models. In the Pythia suite, we observe that larger models outperform smaller counterparts on the same training data.
@@
Developmental alignment of the PLMs on the psychometric intelligence tasks is shown in Figure 4. We make the following key observations:
@@
  • Training shows a region of development: For all four tasks, we see a window of monotonic development in which all models gain the respective cognitive abilities. Variation in the onsets of these windows replicates what is known of cognitive development. For example, children acquire language early (i.e., during the preschool years), whereas the onset of improving fluid reasoning is later, when children enter elementary school, and continues for longer, throughout adolescence. Correspondingly, the models significantly develop linguistic abilities while training on 250 Million to 7 Billion tokens, whereas they acquire fluid reasoning abilities while training on 1 Billion to 20 Billion tokens.
  • After development, training appears to serve an engineering goal: After the window of development, training appears to only serve the engineering goal of loss reduction. This observation is especially pronounced for numeric abilities and conceptual understanding.
  • Assessments for Fluid Reasoning and Linguistic Abilities show significant gains in scaling and greater pre-training: For the Fluid Reasoning and Linguistic Abilities assessments, we see that the alignment score continues to increase as the PLMs are trained on a greater number of tokens. Furthermore, for these abilities, models also show scaling effects, with larger models outperforming smaller ones.
@@ 6 Conclusions
This paper investigates the evidence for psychometric intelligence in PLMs by introducing the novel benchmark ThinkTank. ThinkTank enables comprehensive evaluation of PLMs on their cognitive and developmental alignment. It uses representative assessments of four facets of human intelligence: numeric abilities, linguistic abilities, conceptual understanding, and fluid reasoning. Most significantly, we find a window of monotonic development in which all models improve approximately linearly on the four cognitive abilities. Before that window, we interpret training as endowing “blank slate” models with the requisite structure for rapid learning. Also notable is the finding of PLM scaling effects for the assessments of linguistic abilities and fluid reasoning. We propose evaluation against ThinkTank as a prerequisite before treating PLMs as a model of human cognition and its development.
@@ 7 Limitations
Some limitations of the work are as follows: (1) ThinkTank is an aggregation of psychometric tests for PLMs. The limitations of each individual test are inherited in the benchmark. (2) The alignment scores may be wrongly interpreted when evaluating PLMs with ThinkTank. Alignment scores show the similarity of PLM outputs to human outputs on psychometric tests and indicate that PLMs do not need explicit neural circuitry for these intelligence tests. We do not suggest these models as proxies for humans in any manner and recommend further testing before use. (3) The developmental alignment of the models points towards the acquisition of human-like performance on the four psychometric assessments in ThinkTank in the range of 100 Million to 20 Billion training tokens. This conclusion has two limitations: Pythia is the only suite of models with available intermediate checkpoints and, while unlikely, the observed developmental trajectories might be artifacts of the pre-training order. (4) The psychometric assessments for PLMs are adapted from similar human psychometric tests. Different ways of adaptation may lead to different results. Furthermore, while representative, these assessments are not exhaustive tests of human intelligence. Future work can expand to other tests like spatial and commonsense reasoning. (5) Some open source models like Llama-2 have larger 70 Billion parameter variants but we lack the compute resources to evaluate them. Large open-source models would lead to appropriate comparisons of performance with commercial models like GPT-4. (6) While our work evaluates changes in cognitive alignment with an increase in model size and the number of pre-training tokens, we do not control for different tuning methodologies like instruction tuning and reinforcement learning with human or artificial intelligence feedback.
@@ 8 Ethical Considerations
There are no major risks associated with conducting this research beyond those associated with working with PLMs. There may be risks in misinterpreting the alignment scores when evaluating with ThinkTank. The psychometric analysis of this study is one-way: we look for human performance characteristics and behaviors in PLMs. PLMs are experimental technologies and future work using this research should proceed with caution. Assessment on ThinkTank indicates PLM alignment – or the lack thereof – to human cognitive behavior. Indications of higher human alignment do not indicate an absolute proxy for humans. The goal of ThinkTank is a pre-cursor assessment of PLMs on their ability to act as cognitive models. Therefore, researchers and users should perform more tests before use.
@@ Appendix A Computational Resources
The models are evaluated on Nvidia A100 GPUs with 80 GB RAM. The evaluation in this paper cumulatively takes 1400 GPU hours. We use the provided APIs by OpenAI and Google for models of the GPT-X family and Gemini respectively.
@@ B.1 Numeric abilities: Magnitude comparison effects
Physical quantities in the world are encoded as logarithmically scaled magnitude representations fechnerElementsPsychophysics1860. While the distance and the ratio effects are the biggest indicators of the presence of such log-scaled magnitude representations and the numerical precision in humans, other human effects also explain the mental number line. These effects are as follows:
  • Distance effect (refer to figure 1 (A) top): The greater the distance |x-y| between two numbers (x, y), the faster the comparison in humans MOYER1967.
  • Size effect: Given two comparisons of the same distance (i.e., of the same value for |x - y|), the smaller the numbers, the faster the comparison parkmanTemporalAspectsDigit1971.
  • Ratio effect (refer to figure 1 (A) bottom): The time taken by humans to compare two numbers (x,y) is a decreasing function of the ratio of the larger number over the smaller number $\frac{max(x,y)}{min(x,y)}$ halberdaIndividualDifferencesNonverbal2008.
  • Multidimensional scaling: Along with the three effects, we investigate the consistency of the latent number representations of PLMs with the human MNL using multidimensional scaling mds_BorgGroenen2005; mds_Ding2018. MDS recovers the latent representation from the cosine (dis)similarities between the vector representations of all pairs of numbers (for a given LLM, layer, and number format). This is evaluated by the correlation between the positions of the numbers 1 to 9 in the MDS solution and the expected values (log(1) to log (9)) of the human MNL (refer to the correlation value in table 4).
@@ B.3 Conceptual Understanding
Table 7: Results for the typicality effects using the three methods

### v3

@@ 3.4 Fluid reasoning
Humans can logically parse information and detect patterns in novel stimuli without having to rely on prior experiences or learned information. This ability is called fluid reasoning Cattell (1963).
@@
Amber Liu et al. (2023)
Falcon Almazrouei et al. (2023)
Starling Zhu et al. (2023)
Llama Touvron et al. (2023)
Mistral Jiang et al. (2023)
Qwen Bai et al. (2023)
@@
Biderman et al. (2023)
Gemini Team et al. (2023)
GPT-3.5-Turbo OpenAI (2023b)
GPT 4 OpenAI (2023a)
@@
We focus on the dominant measure of fluid reasoning, the Ravens Progressive Matrices (RPM) test Raven (2003). An example Ravens-like problem is given in Figures 1D and 3. An RPM item consists of a 3x3 matrix of cells with one empty cell. Participants must induce the underlying, abstract patterns that hold across the rows and columns of the matrix, and apply these to infer the image in the empty cell from a given set of options. These images vary in visual attributes like shape and color, along with more abstract qualities. The RPM is the standard measure of fluid reasoning Snow et al. (1984) and is highly correlated with analogical reason Goswami (1986); Webb et al. (2023).
@@
Given the visual nature of the RPM, previous work by Hu et al. (2021, 2023) and Webb et al. (2023) mapped the Raven-10000 dataset to a textual format to facilitate the testing of PLMs. The mapping involves reformulating visual elements into text-based numerical tuples representing attributes like shape, size, and color textually, as illustrated in Figure 3, to form the I-Raven dataset. We use their approach with a focus on the “Center Single Alignment” sub-task, which features a single shape per matrix cell. We differ from their work by evaluating a broader set of models.
@@
Table 3: Performance of Pre-trained Language Models on the tasks. Distance Effect: Averaged $R^{2}$ values of different LLMs when fitting a linear function on the cosine-similarity vs. distance plot. Ratio Effect: Averaged $R^{2}$ values of different LLMs when fitting a negative exponential function on the cosine-similarity vs. ratio plot. Note: Each value is averaged across all three input types and all model layers to produce one generalizable score. Latent Rep: Average Spearman’s Correlation when using the cosine similarity and latent representation-based approach (Note: * refers to the prompting approaches for select models which are gated by APIs, and not the latent representation-based approach), Zero-Shot: Average Spearman’s Correlation when using the zero-shot surprisal values, BLiMP: The Benchmark of Linguistic Minimal Pairs for English, RPM: Raven’s Progressive Matrices
Operationalization: We determine the model’s preferred answer for a problem by comparing the surprisal values of the whole sequence (instruction, question, candidate tuple) for each of the candidate options, i.e. the probability of each completed digit representation of a matrix. For the example given in Figure 3, this would be checking the probability of this sequence (summation of token probabilities) with the correct answer (3, 0.6, 0.8) to the other candidates. A complete list of the prompts used in this paper is given in Appendix B.4.
@@ 4 Models under consideration
We evaluate a wide range of language model families, shown in Table 2. These models are selected based on the following criteria:
Public availability: Open-source models allow us to perform a thorough analysis by accessing the latent representation and the token probability during generation. We follow Holt et al. (2024) while choosing PLMs. Although most models in this study are publicly available and open-source, we use three state-of-art commercial PLMs that are gated behind API calls; GPT-3.5-Turbo (pointing to gpt-3.5-turbo-0613 on the OpenAI platform), GPT-4 (pointing to gpt-4-1106 on the OpenAI platform), and Gemini (also referred to as Gemini-1-Pro at the time of writing). The GPT-$x$ model APIs provide token probabilities of the response, allowing us to calculate surprisal, while Gemini does not.
@@
Availability of intermediate training checkpoints: This allows us to evaluate the effects of pre-training on the model outputs. Together, the availability of multiple model sizes and intermediate training checkpoints allow us to best evaluate the developmental alignment of PLMs. Amber and Pythia’s family of models have available intermediate training checkpoints. While Amber has 360 intermediate checkpoints, the checkpoints are at 4 Billion tokens each and are not at the required granularity.
Pythia Family of models: Pythia Biderman et al. (2023) is one of the first open-source projects with the goal of scientific and transparent model development. It has 8 model sizes ranging from 70 Million to 12 Billion parameters, with each model trained on 286 Billion tokens. The models in the suite are equivalent (in size) to popular decoder architectures like GPT-Neo-(125M, 1.3B, 2.7B) and OPT-(125M, 350M, 1.3B, 2.7B, 6.7B), but with the added benefits of training on a known de-duplicated corpus Gao et al. (2020), using the same training order for each model size, and having 154 intermediate checkpoints to study the learning trajectories of PLMs. Thus, the Pythia suite of models is ideal for studying the cognitive and developmental alignment of PLMs to humans.
All open-source models are obtained from Huggingface Wolf et al. (2020), while the gated models are obtained from their respective platforms through API calls. For each model in the Pythia suite, the following intermediate checkpoints are available: [1, 2, 4, 8, … 512; 1000, 2000, 3000 … 143000 (exponential increase in checkpoint number until the 512th checkpoint and subsequent progression of 1000 steps until the last checkpoint)], with each checkpoint representing 2 Million tokens seen. Overall, we test 1232 intermediate checkpoints of the Pythia suite of models across all the tasks.
@@ 5 Cognitive and developmental alignment of PLMs
The suite of tasks enables comprehensive evaluation of a variety of PLMs on their cognitive alignment to humans across four domains of psychometric intelligence: numeric abilities, linguistic abilities, concept understanding, and fluid reasoning. Table 3.4 highlights the key results of this evaluation. For the evaluation of conceptual understanding in PLMs, we only report the results for the zero-shot surprisal values and latent representations. This is because we see similar results for zero-shot and few-shot surprisal value-based methods (see comprehensive results in Appendix B.3).
@@
  • Linguistic abilities: The accuracy of the PLMs on the BLiMP linguistic acceptability tasks improves upon increasing the number of parameters. Furthermore, we find that all PLMs are substantially more accurate on morphological tasks over syntactic and semantic tasks (Accuracy: semantic $<$ syntax $\ll$ morphology; see Appendix Table 5).
  • Concept understanding: Prompting methods in commercial models perform substantially better than other methods – closeness judgment and surprisal values – on all open-source models. In the Pythia suite, we observe that larger models outperform smaller counterparts on the same training data.
@@
The developmental alignment of the PLMs on the tasks is shown in Figure 4. We make the following key observations:
@@
  • Training shows a region of development: For all four tasks, we see a window of monotonic development, in which all models gain the respective cognitive abilities.
  • After development, training appears to serve an engineering goal: After the window of development, the metric becomes unstable once the phenomena are learned. The training appears to serve the engineering goal of loss reduction Chen et al. (2023). This observation is especially pronounced for numeric abilities and conceptual understanding.
  • Assessments for Fluid Reasoning and Linguistic Abilities show significant gains with scaling and greater pre-training: For the assessments of these abilities, we see that the alignment score continues to increase as the PLMs are trained on a greater number of tokens. (Also, morphological performance develops first followed by syntax and then semantics; see Appendix Figure 7.) Furthermore, for these abilities, models also show scaling effects, with larger models outperforming smaller ones.
  • The relative positions of the windows weakly align with human development: Variation in the onsets of the windows is weakly consistent with what is known of cognitive development. For example, children acquire language early (i.e., during the preschool years), whereas the onset of improving fluid reasoning is later, when children enter elementary school, and continues for longer, throughout adolescence. Correspondingly, the models significantly develop linguistic abilities while training on 250 Million to 7 Billion tokens, whereas they acquire fluid reasoning abilities later, while training on 1 to 20 Billion tokens.
@@ 6 Conclusions
This paper investigates the appropriateness of using PLMs for human cognitive and developmental modeling. It uses representative assessments of four facets of psychometric intelligence: numeric abilities, linguistic abilities, conceptual understanding, and fluid reasoning. Our experiments show that PLMs develop cognitive abilities purely through their experience in the world, indicating that the cognitive abilities we test are acquirable through mere exposure to language distributions and do not necessarily require innate human-like inductive biases. Most significantly, we find a window of monotonic development in which all models improve approximately linearly on the four cognitive abilities. Before that window, we interpret training as endowing “blank slate” models with the requisite structure for rapid learning. Also notable is the finding of PLM scaling effects for the assessments of linguistic abilities and fluid reasoning. We propose evaluation against these tasks as a prerequisite before treating PLMs as models of human cognition and its development.
@@ 7 Limitations
Some limitations of the work are as follows: (1) We use an aggregation of psychometric tests for PLMs. The limitations of each test are inherited in the suite of tasks. (2) The alignment scores may be wrongly interpreted when evaluating PLMs with these tasks. Alignment scores show the similarity of PLM outputs to human outputs on psychometric tests and indicate that PLMs do not need explicit neural circuitry for these intelligence tests. We do not suggest these models as proxies for humans in any manner and recommend further testing before use. (3) The developmental alignment of the models points towards the acquisition of human-like performance on the four psychometric assessments in the range of 100 Million to 20 Billion training tokens. This conclusion has two limitations: Pythia is the only suite of models with available intermediate checkpoints and, while unlikely, the observed developmental trajectories might be artifacts of the pre-training order. (4) The psychometric assessments for PLMs are adapted from similar human psychometric tests. Different ways of adaptation may lead to different results. Furthermore, while representative, these assessments are not exhaustive tests of human intelligence. Future work can expand to other tests like spatial and commonsense reasoning. (5) Some open source models like Llama-2 have larger 70 Billion parameter variants but we lack the compute resources to evaluate them. Large open-source models would lead to appropriate comparisons of performance with commercial models like GPT-4. (6) While our work evaluates changes in cognitive alignment with an increase in model size and the number of pre-training tokens, we do not control for different tuning methodologies like instruction tuning and reinforcement learning with human or artificial intelligence feedback. Accounting for different tuning methods is computationally intensive for the 1200+ model checkpoints across 10 architectures.
@@ 8 Ethical Considerations
All tasks and corresponding datasets have low ethical risks and none expose sensitive information. Additionally, we obtain approval from the authors of each dataset for their use and release. There are no major risks associated with conducting this research beyond those associated with working with PLMs. There may be risks in misinterpreting the alignment scores when evaluating with the tests. The psychometric analysis of this study is one-way: we look for human performance characteristics and behaviors in PLMs. PLMs are experimental technologies and future work using this research should proceed with caution. Assessment of the tasks indicates PLM alignment – or the lack thereof – to human cognitive behavior. Indications of higher human alignment do not indicate an absolute proxy for humans. The goal of tasks in this work is a pre-cursor assessment of PLMs on their ability to act as cognitive models. Therefore, researchers and users should perform more tests before use.
@@ Appendix A Computational Resources
The models are evaluated on Nvidia A100 GPUs with 80 GB RAM. The evaluation in this paper cumulatively takes 1600 GPU hours. We use the provided APIs by OpenAI and Google for models of the GPT-X family and Gemini respectively.
@@ B.1 Numeric abilities: Magnitude comparison effects
Physical quantities in the world are encoded as logarithmically scaled magnitude representations Fechner (1860). While the distance and the ratio effects are the biggest indicators of the presence of such log-scaled magnitude representations and the numerical precision in humans, other human effects also explain the mental number line. These effects are as follows:
  • Distance effect (refer to figure 1 (A) top): The greater the distance |x-y| between two numbers (x, y), the faster the comparison in humans Moyer and Landauer (1967).
  • Size effect: Given two comparisons of the same distance (i.e., of the same value for |x - y|), the smaller the numbers, the faster the comparison Parkman (1971).
  • Ratio effect (refer to figure 1 (A) bottom): The time taken by humans to compare two numbers (x,y) is a decreasing function of the ratio of the larger number over the smaller number $\frac{max(x,y)}{min(x,y)}$ Halberda et al. (2008).
  • Multidimensional scaling: Along with the three effects, we investigate the consistency of the latent number representations of PLMs with the human MNL using multidimensional scaling Borg and Groenen (2005); Ding (2018). MDS recovers the latent representation from the cosine (dis)similarities between the vector representations of all pairs of numbers (for a given LLM, layer, and number format). This is evaluated by the correlation between the positions of the numbers 1 to 9 in the MDS solution and the expected values (log(1) to log (9)) of the human MNL (refer to the correlation value in table 4).
@@ B.2 Linguistic Abilities
Figure 7: Developmental trajectory of the Pythia suite of models on the BLiMP linguistic acceptability tasks.
@@ B.3 Conceptual Understanding
Note: Under the closeness judgment protocol, our experiments fail to match up to the performance of the models used by Vemuri et al. (2024). This is because our choice of open-source models only provides token representations, on which we later perform an aggregation operation. This aggregation operation leads to a loss of information. In contrast, Vemuri et al. (2024) use sentence-transformer models Reimers and Gurevych (2019), which provide singular latent representation for longer text. This variation in experimentation leads to the difference in alignment scores.
@@
Table 7: Results for the typicality effects using the three methods
@@ B.4 Fluid Reasoning
Humans cannot completely operate without relying on prior experience. The pervasive role of prior knowledge in shaping cognition is a foundational tenet of the cognitive revolution. However, “Fluid intelligence” is the ability to solve novel and abstract problems Raven (2003). It is a core cognitive ability, closely related to other domain-general cognitive abilities like working memory, and executive function, both correlationally Conway et al. (2002) and in terms of the underlying neural correlates (i.e., in the prefrontal cortex) Burgess et al. (2011). It is distinguished from crystallized intelligence, which is composed of the domain-specific knowledge and skills one acquires through one’s lifetime Hartshorne and Germine (2015). This distinction is a classic one in psychology Carroll (1993).